### v1

# PCF Learned Sort: a Learning Augmented Sort Algorithm with $\mathcal{O}(n\log\log n)$ Expected Complexity

Abstract
Sorting is one of the most fundamental algorithms in computer science. Recently, Learned Sorts, which use machine learning to improve sorting speed, have attracted attention. While existing studies show that Learned Sort is experimentally faster than classical sorting algorithms, they do not provide theoretical guarantees about its computational complexity. We propose PCF Learned Sort, a theoretically guaranteed Learned Sort algorithm. We prove that the expected complexity of PCF Learned Sort is $\mathcal{O}(n\log\log n)$ under mild assumptions on the data distribution. We also confirm experimentally that PCF Learned Sort has a computational complexity of $\mathcal{O}(n\log\log n)$ on both synthetic and real datasets. This is the first study to theoretically support the experimental success of Learned Sort, and provides evidence for why Learned Sort is fast.
sorting, learning-augmented algorithms, algorithms with predictions
[…]
do
1 Introduction
The sorting algorithm is one of the most important computer algorithms and has been studied for a long time. Recently, a novel sorting method called Learned Sort has been proposed (Kraska et al., 2019). Learned Sort uses a model that predicts the cumulative density function (CDF) to predict the position of each element in the sorted array, and then uses this prediction to quickly sort the array. Experimental results show that Learned Sort is faster than classical sorting algorithms, including highly tuned counting-based sorting algorithms, comparison sorting algorithms, and hybrid sorting algorithms.
On the other hand, there are few theoretical guarantees about the computational complexity of Learned Sort. The first proposed Learned Sort algorithm (Kraska et al., 2019) has a best-case complexity of $\mathcal{O}(n)$, but its expected or worst-case complexity isn’t discussed. The more efficient Learned Sort algorithms proposed later (Kristo et al., 2020, 2021) also have $\mathcal{O}(n)$ best-case complexity, but $\mathcal{O}(n^{2})$ worst-case complexity (or $\mathcal{O}(n\log n)$ with some modifications). The goal of this paper is to provide a Learned Sort that is theoretically guaranteed to be computationally efficient.
We propose PCF Learned Sort, which can sort with expected complexity $\mathcal{O}(n\log\log n)$ under mild assumptions on the data distribution. Furthermore, we prove that PCF Learned Sort can sort with worst-case computational complexity $\mathcal{O}(n\log n)$. Then, we experimentally confirmed that our Learned Sort can sort with the average complexity of $\mathcal{O}(n\log\log n)$ on both synthetic and real datasets.
In Section 2, we first present related work. Then, in Section 3, we introduce our PCF Learned Sort and give the theorems and brief explanations of their proofs that guarantee its computational complexity. Then, in Section 4, we verify our theorems experimentally. We then discuss the limitations and possible future work in Section 5. Finally, in Section 6, we summarize our results and conclusions.
2 Related Work
Our research is in the context of algorithms with predictions, in particular, the sorting algorithms with predictions (Section 2.1). There are two types of sorting algorithms, comparison sorts (Section 2.2) and distribution sorts (Section 2.3), and our proposed method is a distribution sort. However, the idea, implementation, and proof of computational complexity of our method are similar to those of sample sort, which is a type of comparison sort. Furthermore, our proposed algorithm and proof of computational complexity are based on those of Learned Index (Section 2.4).
2.1 Algorithms with predictions
Our research is in the area of algorithms with predictions (Mitzenmacher & Vassilvitskii, 2022). These studies focus on improving the performance (e.g., speed and memory usage) of algorithms for problems such as ski rental (Purohit et al., 2018; Gollapudi & Panigrahi, 2019; Shin et al., 2023), caching (Narayanan et al., 2018; Rohatgi, 2020; Lykouris & Vassilvitskii, 2021; Im et al., 2022), scheduling (Gollapudi & Panigrahi, 2019; Lattanzi et al., 2020; Lassota et al., 2023), and matching (Antoniadis et al., 2023; Dinitz et al., 2021; Sakaue & Oki, 2022) by using predictions from oracles such as machine learning models. The main idea of these novel algorithms is to exploit the characteristics of the data by actively exploiting information not explicitly used by classical algorithms.
In particular, our work is based on research about sorting with predictions. (Kraska et al., 2019) proposed a method for fast sorting that uses a model $\tilde{F}(q)$ trained to approximate the cumulative distribution function (CDF) $F(q)$. This first Learned Sort roughly sorts the array by placing each element $x$ of the array on the $nF(x)$-th entry of the output array. It then sorts the array completely by refining the roughly sorted array using insertion sort. This algorithm has a best case complexity of $\mathcal{O}(n)$, but a worst case complexity of $\mathcal{O}(n^{2})$ because of the exception handling for collisions and the last refining.
Subsequently, implementations have been proposed that improve cache efficiency (Kristo et al., 2020) and are robust against key duplication (Kristo et al., 2021). These also use insertion sorting in the last refinement, resulting in a worst-case complexity of $\mathcal{O}(n^{2})$. The worst-case complexity can be reduced to $\mathcal{O}(n\log n)$ by using a sorting algorithm with small worst-case complexity, such as Introsort (Musser, 1997) or TimSort (McIlroy, 1993). However, this is the same complexity as many effective classical sorting algorithms, and thus does not explain the experimental success of Learned Sort. Some studies have focused on efficiently exploiting the predictions of the oracle, considering the oracle as a black box and not focusing on its inner details (Bai & Coester, 2023). This analysis does not take into account the computational complexity of training and inferring the oracle, so it cannot be directly applied to guarantee the overall computational complexity of sorting.
2.2 Comparison sorts
Sorting algorithms that use comparisons between keys and require no other information about the keys are called comparison sorts. It can be proved that the worst complexity of a comparison sort is at least $\mathcal{O}(n\log n)$. Commonly used comparison sorting algorithms include quick sort, heap sort, merge sort, and insertion sort. The GNU Standard Template Library in C++ uses Introsort (Musser, 1997), an algorithm that combines quick sort, heap sort, and insertion sort. Java (Java, 2023) and Python up to version 3.10 (Peters, 2002) use TimSort (McIlroy, 1993), an improved version of merge sort. Python 3.11 and later use Powersort (Munro & Wild, 2018), a merge sort that determines a near-optimal merge order.
Sample sort (Frazer & McKellar, 1970) is an extension of quick sort that uses multiple pivots, whereas quick sort uses only one pivot. Sample sort samples a small number of keys from the array, determines multiple pivots, and uses them to partition the array into multiple buckets. The partitioning is repeated recursively until the array is sufficiently small. The in-place parallel superscalar sample sort (Axtmann et al., 2022) is one of the most efficient sample sort implementation. They theoretically guarantee its computational and cache efficiency by proving a theorem about the probability that the pivots partition the array (nearly) equally.
2.3 Distribution sorts
Distribution sorts use information other than key comparison. Radix sort and counting sort are the most common types of distribution sorts. Radix sort uses counting sort as a subroutine for each digit. When the number of digits in the array element is $w$, the computational complexity of radix sort is $\mathcal{O}(wn)$. Thus, radix sort is particularly effective when the number of digits is small. There are several variants of radix sort, such as SkaSort (Skarupke, 2016), which is optimized using a technique called “American flag sort,” and RegionSort (Obeya et al., 2019), which enables efficient parallelization by modeling and resolving dependencies among the element swaps.
2.4 Learned Index
Kraska et al. showed that index data structures such as B-trees and Bloom filters can be made faster or more memory efficient by combining them with machine learning models, and named such novel data structures Learned Index (Kraska et al., 2018). Since then, various learning augmented B-trees (Wang et al., 2020; Kipf et al., 2020) and learning augmented Bloom filters (Mitzenmacher, 2018; Vaidya et al., 2021; Sato & Matsui, 2023) have been proposed. There are several researches on learning augmented B-trees whose performance is theoretically guaranteed. PGM-index (Ferragina & Vinciguerra, 2020) is a learning augmented B-tree that is guaranteed to have the same worst-case query complexity as the classical B-tree, i.e., $\mathcal{O}(\log n)$. (Zeighami & Shahabi, 2023) proposed a learning augmented B-tree with an expected query complexity of $\mathcal{O}(\log\log n)$ under mild assumptions on the distribution.
3 Methods
Figure 1: PCF Learned Sort: First, the input array is partitioned into $\gamma+1$ buckets using a CDF model-based method. Buckets larger than $\delta$ or smaller than $\tau$ are sorted with a “standard” sorting algorithm of complexity $\mathcal{O}(n\log n)$ (e.g. IntroSort). Otherwise, the recursive model-based bucketing is repeated. Finally, the sorted arrays are concatenated. The CDF model used for bucketing is a Piecewise Constant Function (PCF). The function is constant within each interval and the interval widths are constant.
Algorithm 1 The Learned Sort algorithm
[…]
This section describes our PCF Learned Sort. First, in Section 3.1, we give an overview of the algorithm and the key theorems used to guarantee the expected and worst-case computational complexity. Next, in Section 3.2, we introduce the specific algorithm for PCF Learned Sort and give a guarantee on its computational complexity.
3.1 Method Overview
Our algorithm basically repeats recursive model-based bucketing until the array is small enough or the bucketing “fails.” In model-based bucketing, a CDF model is trained and then used to partition an array of length $n$ into approximately $\gamma$ buckets, satisfying the following two conditions: the elements in $i$-th bucket are smaller than any elements in $(i+1)$-th bucket, and the buckets have approximately the same size with high probability. If we set $\gamma=n^{c}$ (where $10<c<$ 1 is a constant) and the partitioning into buckets is done well, the size of each bucket will be approximately $n^{1-c}$, so the size of the bucket in the $k$-th recurrence depth will be approximately $n^{(1-c)^{k}}$. Here,
$k\geq-\frac{\log\log n}{\log(1-c)}\Rightarrow n^{(1-c)^{k}}$ $\leq n^{(1-c)^{-\frac{\log\log n}{\log(1-c)}}}$ (1)
$=n^{\frac{1}{\log n}}$ (2)
$=e,$ (3)
where $\log$ is the natural logarithm, so the base is $e$. That is, with a recursion depth of $k=\Omega(\log\log n)$, the size of bucket is small enough, i.e., the max recursion depth is $\mathcal{O}(\log\log n)$. Thus, if the expected computational complexity of the partition is $\mathcal{O}(n)$ and the probability that the bucketing “fails” is sufficiently small, the total computational complexity is $\mathcal{O}(n\log\log n)$. Under mild assumptions on the distribution, we can prove that this fast and accurate partition can be achieved by using the Piecewise Constant Function (PCF) as the CDF model.
Let $\mathcal{D}~{}(\subset{\mathbb{R}})$ be the range of possible values of the input array and $\bm{x}~{}(\in\mathcal{D}^{n})$ be the input array. If the length of the input array, denoted as $|\bm{x}|$, is less than $\tau$, our algorithm sorts the input array using a “standard” sorting method with $\mathcal{O}(n\log n)$ complexity, such as IntroSort, where $\tau$ is a predetermined constant. Otherwise, if $|\bm{x}|\geq\tau$, model-based bucketing is performed.
The model-based bucketing method $\mathcal{M}$ takes an input array $\bm{x}$ and partitions it into several buckets. First, it determines the parameter $\gamma~{}(\in\mathbb{N})$, which determines the number of buckets. $\gamma$ is determined as a function of the array length $n$. The number of buckets returned by $\mathcal{M}$ is $\gamma+1$. Next, all or some elements of $\bm{x}$ are used to train the CDF model $\tilde{F}\colon\mathcal{D}\mapsto[0,1]$. The $\tilde{F}(q)$ is trained to predict the empirical cumulative distribution function of $\bm{x}$ defined as:
$F_{\bm{x}}(q)\coloneqq\frac{|\{i\in\{1,\dots,n\}\mid x_{i}\leq q\}|}{|\bm{x}|}.$ (4)
The specific model and training method is explained in Section 3.2. Finally, the CDF model is used to partition the input array $\bm{x}$ into $\gamma+1$ buckets. All $\gamma+1$ buckets, $\{\bm{c}_{j}\}_{j=1}^{\gamma+1}$, are initialized to be empty, and then for each $i\in\{1,\dots,n\}$, $x_{i}$ is appended to $\bm{c}_{\left\lfloor\tilde{F}(x_{i})\gamma\right\rfloor+1}$. This is based on the intuition that the number of elements less than or equal to $x_{i}$ in the array $\bm{x}$ (i.e., |$|\{j\in\{1,\dots,n\}\mid x_{j}\leq x_{i}\}|$) is approximately equal to $n\tilde{F}(x_{i})$.
The CDF model $\tilde{F}$ must be a non-decreasing function to guarantee the computational complexity of our algorithm. In this case, the bucket with the larger ID gets the larger value, i.e.,
$p\in\bm{c}_{j},q\in\bm{c}_{k},1\leq j<k\leq\gamma+1\Rightarrow p<q.$ (5)
This means that each bucket is responsible for a disjoint and continuous interval. Let $t_{j}=\min_{x\in\bm{c}_{j}}x~{}(j=1,\dots,\gamma+1),~{}t_{\gamma+2}=\infty$, then the $j$-th bucket $\bm{c}_{j}$ ($j=1,\dots,\gamma+1$) is responsible for a continuous interval $\mathcal{I}_{j}\coloneqq[t_{j},t_{j+1})$.
After model-based bucketing, our algorithm determines for each bucket whether the bucketing was “successful” or not. For each $j=1,\dots,\gamma+1$, we check whether the size of bucket $\bm{c}_{j}$ is less than $\delta$, where $\delta$ is an integer determined by $n$. If $|\bm{c}_{j}|\geq\delta$ (which means the bucketing “fails”), the bucket is sorted using the “standard” sorting method (e.g., IntroSort). If $|\bm{c}_{j}|<\delta$ (which means the bucketing “succeeds”), the bucket is sorted by recursively calling our Learned Sort algorithm. Note that the parameters such as $\gamma$ and $\delta$ are redetermined according to the size of the bucket (i.e., the input array in the next recursion step), and the CDF model is retrained for each bucket. After each bucket is sorted, the buckets are concatenated. The input array $\bm{x}$ is sorted by the above procedure. The basic idea of our algorithm is visualized in Figure 1, and the pseudocode is given in Algorithm 1.
The following is a lemma about the worst-case complexity of our Learnd Sort as defined above.
Lemma 3.1.
Assume that there exists an model-based bucketing algorithm $\mathcal{M}$ such that $\mathcal{M}$ can perform bucketing (including model training and inferences) an array of length $n$ into $\gamma+1=\mathcal{O}(n)$ buckets with a worst-case complexity of $\mathcal{O}(n)$. Also, assume that the “standard” sort algorithm used for the too large or too small buckets has a worst-case complexity of $\mathcal{O}(n\log n)$. The worst-case complexity of our Learned Sort with such $\mathcal{M}$ and $\delta=\lfloor n^{d}\rfloor$ (where $d$ is a constant satisfying $10<d<$ 1) is $\mathcal{O}(n\log n)$.
This lemma can be intuitively shown from the following two points: (i) the maximum recursion depth is $\mathcal{O}(\log\log n)$, and (ii) each element of the input array $\bm{x}$ undergoes several bucketing and only one “standard” sort. (i) can be shown from the fact that the size of the bucket in the $i$-th recursion depth is less than $n^{d^{i}}$, and (ii) is evident from the algorithm’s design since the buckets sorted by “standard” sort are now left only to be concatenated. The exact proof is given in Section A.1.
Next, we introduce an important lemma about the expected complexity of our Learned Sort. The following assumption is necessary to guarantee the expected complexity.
Assumption 3.2.
The input array $\bm{x}\in\mathcal{D}^{n}$ is formed by independent sampling according to a probability density distribution $f(x)\colon\mathcal{D}\mapsto\mathbb{R}_{\geq 0}$.
We define $f_{\mathcal{I}}(x)\colon\mathcal{I}\mapsto\mathbb{R}_{\geq 0}$ to be the conditional probability density distribution of $f(x)$ under the condition that $x\in\mathcal{I}$ for a interval $\mathcal{I}\subset\mathcal{D}$, i.e.,
[…]
The expected computational complexity of our proposed Learned Sort is guaranteed by the following lemma.
Lemma 3.3.
Let $\bm{x}_{\mathcal{I}}~{}(\in\mathcal{I}^{n})$ be the array formed by sampling $n$ times independently according to $f_{\mathcal{I}}(x)$. Assume that there exist a model-based bucketing algorithm $\mathcal{M}$ and a constant $d~{}(\in(0,1))$ that satisfy the following for any interval $\mathcal{I}~{}(\subset\mathcal{D})$:
  • $\mathcal{M}$ can perform bucketing (including model training and inferences) an array of length $n$, with an expected complexity of $\mathcal{O}(n)$. That is, $\mathcal{M}$ can take the array $\bm{x}_{\mathcal{I}}~{}(\in\mathcal{I}^{n})$ as input and divide it into $\gamma+1$ buckets, $\{\bm{c}_{j}\}_{j=1}^{j=\gamma+1}$ satisfying Equation 5, with an expected complexity of $\mathcal{O}(n)$.
  • $\gamma+1=\mathcal{O}(n)$.
  • $\Pr[\exists j,|\bm{c}_{j}|\geq\lfloor n^{d}\rfloor]=\mathcal{O}\left(\frac{1}{\log n}\right)$.
The expected complexity of our Learned Sort with such $\mathcal{M}$ and $\delta=\lfloor n^{d}\rfloor$ is $\mathcal{O}(n\log\log n)$.
This lemma can be proved intuitively by the following two points: (i) the maximum recursion depth is $\mathcal{O}(\log\log n)$, and (ii) the expected total computational complexity from the $i$-th to the $(i+1)$-th recursion depth is $\mathcal{O}(n)$. (i) is the same as in the explanation of the proof of Lemma 3.1. (ii) can be shown from the fact that the expected computational complexity from the $i$-th to the $(i+1)$-th recursion depth is $\mathcal{O}(n\log n)$ with probability $\mathcal{O}(\frac{1}{\log n})$, and $\mathcal{O}(n)$ in other cases. See Section A.2 for the exact proof.
Note that the assumption of Lemma 3.3 includes “$\mathcal{M}$ works well with high probability for any $\mathcal{I}$.” This is because our Learned Sort algorithm recursively repeats the model-based bucketing. The range of elements in the bucket, i.e., the input array in the next recursion step, can be any interval $\mathcal{I}~{}(\subset\mathcal{D})$.
3.2 PCF Learned Sort
We propose PCF Learned Sort as an implementation that satisfies the assumptions of Lemma 3.1 and Lemma 3.3, and thus has a guarantee that the worst-case complexity is $\mathcal{O}(n\log n)$ and the expected complexity is $\mathcal{O}(n\log\log n)$. PCF Learned Sort approximates the CDF by a Piecewise Constant Function (PCF). Piecewise constant function is a function that has intervals of equal width and outputs the same value in each interval (the right side of Figure 1). The study that develops a Learned Index with a theoretical guarantee on its complexity (Zeighami & Shahabi, 2023) also used PCF as a CDF model.
The model-based bucketing method in PCF Learned Sort $\mathcal{M}_{\mathrm{PCF}}$ trains the CDF model $\tilde{F}$ in the following way. First, the parameters $\alpha\in\{1,\dots,n\}$ and $\beta\in\mathbb{N}$ are determined by $n$, the length of the input array. The parameter $\alpha$ is the number of samples to train the CDF model, and $\beta$ is the number of intervals in the PCF. Next, the piecewise constant function is trained by counting the number of samples in each interval. Using $x_{\mathrm{min}}=\min_{i}{x_{i}}$ and $x_{\mathrm{max}}=\max_{i}{x_{i}}$, let $i(x)$ be a function defined as follows:
$i(x)=\left\lfloor\frac{x-x_{\mathrm{min}}}{x_{\mathrm{max}}-x_{\mathrm{min}}}\beta\right\rfloor+1.$ (7)
$\alpha$ samples are taken at random from $\bm{x}$ to form $\bm{a}~{}(\in\mathcal{D}^{\alpha})$, and then $i(x)$ is used to form the array $\bm{b}~{}(\in{\mathbb{Z}_{\geq 0}}^{\beta+1})$ defined as follows:
$b_{i}=\left|\left\{j\in\{1,\dots,\alpha\}\mid i(a_{j})\leq i\right\}\right|.$ (8)
This counting is corresponding to the training of the PCF. Note that $\bm{b}$ is an non-decreasing non-negative array and $b_{\beta+1}=\alpha$, i.e., $0\leq b_{1}\leq b_{2}\leq\dots\leq b_{\beta+1}=\alpha$.
The inference for the CDF model $\tilde{F}(x)$ is performed using $i(x)$, $\bm{b}$, and the following equation:
$\tilde{F}(x)=\frac{b_{i(x)}}{\alpha}.$ (9)
Since $i(x)$ is a non-decreasing function and $\bm{b}$ is also a non-decreasing array, $\tilde{F}(x)$ is a non-decreasing function. Also, $0\leq\tilde{F}(x)\leq 1$ because $0\leq b_{i}\leq\alpha$ for every $i$.
The following is a lemma to bound the probability that $\mathcal{M}_{\mathrm{PCF}}$ will “fail” bucketing. This lemma is important to guarantee the expected computational complexity of PCF Learned Sort.
Lemma 3.4.
Let $\sigma_{1}$ and $\sigma_{2}$ be respectively the lower and upper bounds of the probability density distribution $f(x)$ in $\mathcal{D}$, and assume that $0<\sigma_{1}\leq\sigma_{2}<\infty$. That is, $x\in\mathcal{D}\Rightarrow\sigma_{1}\leq f(x)\leq\sigma_{2}$.
Then, in model-based bucketing of $\bm{x}_{\mathcal{I}}~{}(\in\mathcal{I}^{n})$ to $\{\bm{c}_{j}\}_{j=1}^{\gamma+1}$ using $\mathcal{M}_{\mathrm{PCF}}$, the following holds for any interval $\mathcal{I}~{}(\subset\mathcal{D})$:
$\begin{split}&K\geq 1\\ \Rightarrow&\Pr[\exists j,|\bm{c}_{j}|>\delta]\leq\frac{2n}{\delta}\exp\left\{-\frac{\alpha K}{2\gamma}\left(1-\frac{1}{K}\right)^{2}\right\},\end{split}$ (10)
where
$K\coloneqq\frac{\gamma\delta}{2n}-\frac{2\sigma_{2}\gamma}{\sigma_{1}\beta}.$ (11)
The proof of this lemma is based on and combines proofs from two existing studies. The first is Lemma 5.2. from a study of IPS4(Axtmann et al., 2022), an efficient sample sort. This lemma guarantees the probability of a “successful recursion step” when selecting pivots from samples and using them to perform a partition. Although the definition of “successful recursion step” in this paper is similar to our definition of “successful bucketing” (more precisely, in the special case of $\delta=3n/\gamma$), this lemma is for the method that do not use the CDF model, so the proof cannot be applied directly to our case. Another proof we refer to is the proof of Lemma 4.5. from a study that addressed the computational complexity guarantee of the Learned Index (Zeighami & Shahabi, 2023). This lemma provides a probabilistic guarantee for the error between the output of the PCF and the empirical CDF. Some modifications are required to adapt it to the context of sorting and to attribute it to the probability of bucketing failure, i.e., $\Pr[\exists j,|\bm{c}_{j}|>\delta]$. By appropriately combining the proofs of these two lemmas, Lemma 3.4 can be proved. The exact proof is given in Section A.3.
Using Lemma 3.1, Lemma 3.3 and Lemma 3.4, we can prove the following theorems.
Theorem 3.5.
If we use $\mathcal{M}_{\mathrm{PCF}}$ as the bucketing method and set $\alpha=\beta=\gamma=\delta=\lfloor n^{3/4}\rfloor$, the worst-case complexity of PCF Learned Sort is $\mathcal{O}(n\log n)$.
Proof.
When $\alpha=\beta=\gamma=\lfloor n^{3/4}\rfloor$, the computational complexity for model-based bucketing is $\mathcal{O}(n)$. This is because (i) the PCF is trained in $\mathcal{O}(\alpha+\beta)=\mathcal{O}(n^{3/4})$, and (ii) the total complexity of inference for $n$ elements is $\mathcal{O}(n)$, since the inference is performed in $\mathcal{O}(1)$ per element. Therefore, from Lemma 3.1, the worst-case complexity of PCF Learned Sort is $\mathcal{O}(n\log n)$. ∎
Theorem 3.6.
Let $\sigma_{1}$ and $\sigma_{2}$ be the lower and upper bounds, respectively, of the probability density distribution $f(x)$ in $\mathcal{D}$, and assume that $0<\sigma_{1}\leq\sigma_{2}<\infty$.
Then, if we use $\mathcal{M}_{\mathrm{PCF}}$ as the bucketing method and set $\alpha=\beta=\gamma=\delta=\lfloor n^{3/4}\rfloor$, the expected complexity of PCF Learned Sort is $\mathcal{O}(n\log\log n)$.
Proof.
When $\alpha=\beta=\gamma=\lfloor n^{3/4}\rfloor$, the computational complexity for model-based bucketing is $\mathcal{O}(n)$.
Since $K=\Omega(\sqrt{n})$ when $\alpha=\beta=\gamma=\delta=\lfloor n^{3/4}\rfloor$, $K\geq 1$ for sufficiently large $n$, and
$\begin{split}\frac{2n}{\delta}\exp\left\{-\frac{\alpha K}{2\gamma}\left(1-\frac{1}{K}\right)^{2}\right\}&=\mathcal{O}(n^{\frac{1}{4}}\exp(-\sqrt{n}))\\ &\leq\mathcal{O}\left(\frac{1}{\log n}\right).\end{split}$ (12)
Therefore, from Lemma 3.3, the expected computational complexity of PCF Learned Sort is $\mathcal{O}(n\log\log n)$. ∎
Note that the exact value of $\sigma_{1}$ and $\sigma_{2}$ is not required to run PCF Learned Sort, since the parameters for this algorithm, i.e., $\alpha$, $\beta$, $\gamma$ and $\delta$, are determined without the knowledge. In other words, PCF Learned Sort can sort in expected $\mathcal{O}(n\log\log n)$ complexity as long as $0<\sigma_{1}\leq f(x)\leq\sigma_{2}<\infty$, even if it does not know the exact value of $\sigma_{1}$ and $\sigma_{2}$. If $\sigma_{1}$ and $\sigma_{2}$ do not satisfy the assumption of Theorem 3.6, i.e., $\sigma_{1}=0$ or $\sigma_{2}=\infty$, then the expected complexity of PCF Learned Sort is increased to $\mathcal{O}(n\log n)$, but from Theorem 3.5, it cannot be greater than $\mathcal{O}(n\log n)$.
4 Experiments
In this section, we confirm our theorems experimentally. First, we confirm that the computational complexity of PCF Learned Sort is $\mathcal{O}(n\log\log n)$ for both synthetic and real data. Then, we conduct experiments with various parameter settings and experimentally confirm Lemma 3.4, a lemma that bounds the probability of bucketing failure and is an important lemma to guarantee the expected computational complexity of PCF Learned Sort.
4.1 Results on Synthetic Datasets
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 2: Number of operations to sort the synthetic array. Below each graph is a histogram that visualizes the distribution of each dataset.
We experimented with datasets created from the following four types of synthetic distributions: uniform distribution ($\mathsf{min}=0,\mathsf{max}=1$), normal distribution ($\mu=0,\sigma=1$), exponential distribution ($\lambda=1$), lognormal distribution ($\mu=0,\sigma=1$). We took $n=10^{3},\dots,10^{8}$ samples from each distribution independently, and meticulously counted the total number of basic operations to sort the sampled array. Here, the basic operations consists of four arithmetic operations, powers, comparisons, logical operations, assignments, and memory access. We chose this metric, which counts basic operations, to mitigate the environmental dependencies observed in other metrics, such as CPU instructions and CPU time, which are heavily influenced by compiler optimizations and the underlying hardware (this is the same idea as the metric selection in the experiment of  (Zeighami & Shahabi, 2023)). The parameters of PCF Learned Sort are set as in Theorem 3.5 and Theorem 3.6, $\alpha=\beta=\gamma=\delta=\lfloor n^{3/4}\rfloor$. As the “standard” sort algorithm used in PCF Learned Sort, we used quick sort, which is known to have an expected computational complexity of $\mathcal{O}(n\log n)$. We then compared the number of operations between PCF Learned Sort and the (plain) quick sort.
Figure 2 shows the number of operations divided by the length of input array, $n$. It shows the mean and standard deviation of the $10$ measurements for each condition. We see that our PCF Learned Sort has up to 2.8 times fewer operations than quick sort. Also, while the graph of quick sort is almost linear, the graph of PCF Learned Sort is almost flat. This suggests that PCF Learned Sort has a computational complexity much smaller than $\mathcal{O}(n\log n)$.
4.2 Results on Real Datasets
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 3: Number of operations to sort the real data array. Below each graph is a histogram that visualizes the distribution of each dataset.
We used four real datasets, NYC, Wiki, OSM, and Books, which are described below. NYC: pick-up datetimes in the yellow taxi trip records (Kristo, 2021). Wiki: Wikipedia article edit timestamps (Marcus et al., 2020). OSM: uniformly sampled OpenStreetMap locations represented as Google S2 CellIds (Marcus et al., 2020). Books: book sale popularity data from Amazon (Marcus et al., 2020). For each dataset, we randomly sample $n=10^{3},\dots,10^{8}$ elements, shuffle the sampled elements, and then use it as an input array to examine the performance of the algorithms. As in the case of the synthetic dataset, $\alpha=\beta=\gamma=\delta=\lfloor n^{3/4}\rfloor$ and quick sort is used as “standard” sort.
Figure 3 shows plots of the number of operations divided by the length of input array, $n$. Again, the mean and standard deviation of $10$ measurements are shown. We see that our PCF Learned Sort has up to 2.7 times fewer operations than quick sort (as in the case of the synthetic dataset), and that PCF Learned Sort has much less computational complexity than $\mathcal{O}(n\log n)$.
We can also see that the number of operations required for the OSM dataset is about $1.25$ times larger than that for the other datasets. Although not mentioned explicitly, the same tendency is observed in the experiments of existing Learned Sort studies (Kristo et al., 2020, 2021). Our theoretical analysis can explain this phenomenon from the perspective of $\sigma_{2}/\sigma_{1}$, in other words, the strength of the distributional bias. From Equation 11 and Equation 10, the larger $\sigma_{2}/\sigma_{1}$ is, the smaller $K$ is, and the larger (the upper bound of) the probability of bucketing failure. Therefore, the larger $\sigma_{2}/\sigma_{1}$ is, the more the computational complexity will move away from $\mathcal{O}(n\log\log n)$ and closer to the worst-case computational complexity, $\mathcal{O}(n\log n)$. In fact, the histograms of each dataset show that the OSM dataset has sharp peaks and many intervals with almost no data, indicating that it has a distribution with large $\sigma_{2}/\sigma_{1}$.
4.3 Confirmation of Lemma 3.4
Figure 4: Heatmap showing the experimental frequency of bucketing failure, i.e., $\exists j,|\bm{c}_{j}|>\delta$. The variables $a,b,c,d$, except those on the x- and y-axes, were set to $0.75$. The white dotted line represents the parameters that make the right side of Equation 10 equal to 0.5.
Lemma 3.4 bounds the probability that a bucket of size greater than $\delta$ exists. This is an important lemma that allows us to guarantee the expected computational complexity of PCF Learned Sort. Here, we experimentally confirm that this upper bound is appropriate.
We have experimented with $\alpha=\lfloor n^{a}\rfloor,\beta=\lfloor n^{b}\rfloor,\gamma=\lfloor n^{c}\rfloor,\delta=\lfloor n^{d}\rfloor$, varying $a,b,c,d$ from 0.05 to 0.95 at 0.05 intervals. For each $a,b,c,d$ setting, the following was repeated $100$ times: we took $n=10^{6}$ elements from the uniform distribution to form the input array and divided the array into $\gamma+1$ buckets by $\mathcal{M}_{\mathrm{PCF}}$, and checked whether or not $\exists j,|\bm{c}_{j}|>\delta$. Thus, for each $a,b,c,d\in\{0.05,0.10,\dots,0.95\}$, we obtained the experimental frequency at which bucketing “fails.”
Heat maps in Figure 4 show the experimental frequency of bucketing failures when two of the $a,b,c,d$ parameters are fixed and the other two parameters are varied. The value of the two fixed variables are set to 0.75, e.g., in the upper left heap map of Figure 4 (horizontal axis is $a$ and vertical axis is $b$), $c=d=0.75$. The white dotted line represents the parameter so that the right side of Equation 10 is 0.5. That is, Lemma 3.4 asserts that “in the region upper right of the white dotted line, the probability of bucketing failure is less than 0.5.”
We can see that the white dotted line is close to (or slightly to the upper right of) the actual bound of whether bucketing “succeeds” or “fails” more often. In other words, we can see that the theoretical upper bound from Lemma 3.4 agrees well (to some extent) with the actual probability. We can also confirm that, as Lemma 3.4 claims, the probability of bucketing failure is indeed small in the region upper right of the white line.
5 Limitation and Future Work
We proposed PCF Learned Sort, whose expected computational complexity can be proved to be $\mathcal{O}(n\log\log n)$ under the assumption that there exist $0<\sigma_{1}$ and $\sigma_{2}<\infty$ satisfying $\sigma_{1}<f(x)<\sigma_{2}$. Thus, our proof is not applicable to the case of a distribution where there exists $x\in\mathcal{D}$ such that $f(x)=0$ or $f(x)=\infty$. The development of a theory and/or algorithm that is applicable for a wider class of distributions is a future work.
Other future work may include incorporating a more advanced CDF approximation method with a theoretical guarantee into our Learned Sort algorithm. Using a different CDF model and a bucketing algorithm that satisfies the assumptions in Lemma 3.3, we may be able to develop an even more efficient sorting algorithm. In other words, a more accurate and faster CDF approximation method can lead to an improvement of the Learned Sort.
6 Conclusion
We proposed PCF Learned Sort and proved theoretically that its worst-case computational complexity is $\mathcal{O}(n\log n)$ (without assumptions) and its expected computational complexity is $\mathcal{O}(n\log\log n)$ under mild assumptions on the distribution. We then confirm this computational complexity experimentally on both synthetic and real data. Furthermore, we provide a theoretical and intuitive explanation for the slow speed of Learned Sort on some datasets. This is the first study to theoretically support the experimental success of Learned Sort and provides insight into why and when Learned Sort is fast.
Impact Statements
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Proofs
In this Appendix, we give the proofs omitted in the main paper. In Section A.1, we give the proof of Lemma 3.1, which is important for proving the worst-case complexity of PCF Learned Sort. Section A.2 and Section A.3 give proofs of Lemma 3.3 and Lemma 3.4, respectively, which are important for proving the expected computational complexity of PCF Learned Sort.
A.1 Proof of Lemma 3.1
Proof.
Let $P(n)$ be the worst-case complexity of our Learned Sort when using the model-based bucketing algorithm $\mathcal{M}$ as assumed in Lemma 3.1 and $\delta=\lfloor n^{d}\rfloor$. Let $S(n)$ be the worst-case complexity of the “standard” sort algorithm and $R(n)$ be the worst-case complexity of model-based bucketing (including model training and inferences). Since $S(n)=\mathcal{O}(n\log n)$ and the “standard” sorting algorithm terminates after a finite number of operations, $R(n)=\mathcal{O}(n)$, and $\gamma+1=\mathcal{O}(n)$,
$\exists~{}C_{1},l_{1}~{}(>0),$ $~{}~{}n\geq 0\Rightarrow S(n)\leq C_{1}+l_{1}n\log n.$ (13)
$\exists~{}n_{2},l_{2}~{}(>0),$ $~{}~{}n\geq n_{2}\Rightarrow R(n)\leq l_{2}n.$ (14)
$\exists~{}n_{3},l_{3}~{}(>0),$ $~{}~{}n\geq n_{3}\Rightarrow\gamma+1\leq l_{3}n.$ (15)
First, for $n<\max(n_{2},n_{3},\tau)\eqcolon n_{0}$, there exists a constant $C~{}(>0)$ such that $P(n)\leq C$. That is, for $n<n_{0}$, our Learned Sort terminates in a finite number of operations. This is because, since $\delta<n$, the bucket will either be smaller than the original array length $n$, or the bucket will be immediately sorted by the “standard” sort algorithm.
Next, assume that there exists a constant $C~{}(>0)$ and $l~{}(>0)$ such that $P(n)\leq C+ln\log n$ for all $n<k$, where $k$ is an integer such that $k\geq n_{0}$. Let $\mathcal{S}_{\gamma,k}$ be the set consisting of all $(\gamma+1)$-dimensional vectors of positive integers whose sum is $k$, i.e., $\mathcal{S}_{\gamma,k}\coloneq\left\{\bm{s}\in\mathbb{Z}_{\geq 0}^{\gamma}\mid\sum_{i=1}^{\gamma+1}s_{i}=k\right\}$. Then,
$\begin{split}P(k)&\leq R(k)+\max_{\bm{s}\in\mathcal{S}}\sum_{i=1}^{\gamma+1}\left\{\mathbbm{1}[s_{i}\geq\lfloor k^{d}\rfloor]\cdot S(s_{i})+\mathbbm{1}[s_{i}<\lfloor k^{d}\rfloor]\cdot P(s_{i})\right\}\\ &\leq l_{2}k+\max_{\bm{s}\in\mathcal{S}}\sum_{i=1}^{\gamma+1}\left\{\mathbbm{1}[s_{i}\geq\lfloor k^{d}\rfloor]\cdot\left(C_{1}+l_{1}s_{i}\log s_{i}\right)+\mathbbm{1}[s_{i}<\lfloor k^{d}\rfloor]\cdot\left(C+ls_{i}\log s_{i}\right)\right\}\\ &\leq l_{2}k+\max_{\bm{s}\in\mathcal{S}}\sum_{i=1}^{\gamma+1}\left\{\left(C_{1}+l_{1}s_{i}\log s_{i}\right)+\left(C+ls_{i}\log k^{d}\right)\right\}\\ &\leq l_{2}k+C_{1}(\gamma+1)+l_{1}k\log k+C(\gamma+1)+lk\log k^{d}\\ &\leq l_{2}k+C_{1}l_{3}k+l_{1}k\log k+Cl_{3}k+lkd\log k\\ &\leq\left\{l_{2}+C_{1}l_{3}+l_{1}+Cl_{3}-l(1-d)\right\}k\log k+(C+lk\log k).\end{split}$ (16)
Therefore, if we take $l$ such that
$\frac{l_{2}+C_{1}l_{3}+l_{1}+Cl_{3}}{1-d}\leq l,$ (17)
then $P(k)\leq C+lk\log k$ (note that the left side of the following equation is a constant independent of $k$).
Hence, by mathematical induction, it is proved that there exists a constant $C~{}(>0)$ and $l~{}(>0)$ such that $P(n)\leq C+ln\log n$ for all $n\in\mathbb{N}$. ∎
A.2 Proof of Lemma 3.3
Proof.
The proof approach is the same as in Lemma 3.1, but in Lemma 3.3 the “expected” computational complexity is bounded. The following two randomnesses are considered to compute the “expected” computational complexity: (i) the randomness with which $n$ elements are independently sampled according to the probability density function $f(x)$ in the process of forming the input array $\bm{x}$, and (ii) the randomness of the PCF Learned Sort algorithm sampling $\alpha$ elements from the input array $\bm{x}$ for training the PCF.
Let $T(n)$ be the expected complexity of our Learned Sort when using the model-based bucketing algorithm $\mathcal{M}$ as assumed in Lemma 3.3 and $\delta=\lfloor n^{d}\rfloor$. Let $S(n)$ be the expected complexity of the “standard” sort algorithm and $R(n)$ be the expected complexity of model-based bucketing (including model training and inferences). Since $S(n)=\mathcal{O}(n\log n)$ and the “standard” sorting algorithm terminates after a finite number of operations, $R(n)=\mathcal{O}(n)$, $\Pr[\exists j,|\bm{c}_{j}|\geq\lfloor n^{d}\rfloor]=\mathcal{O}(1/\log n)$, and $\gamma+1=\mathcal{O}(n)$,
$\exists~{}C_{1},l_{1}~{}(>0),$ $~{}~{}n\geq 0\Rightarrow S(n)\leq C_{1}+l_{1}n\log n,$ (18)
$\exists~{}n_{2},l_{2}~{}(>0),$ $~{}~{}n\geq n_{2}\Rightarrow R(n)\leq l_{2}n,$ (19)
$\exists~{}n_{3},l_{3}~{}(>0),$ $~{}~{}n\geq n_{3}\Rightarrow\Pr[\exists j,|\bm{c}_{j}|\geq\lfloor n^{d}\rfloor]\leq\frac{l_{3}}{\log n},$ (20)
$\exists~{}n_{4},l_{4}~{}(>0),$ $~{}~{}n\geq n_{4}\Rightarrow\gamma+1\leq l_{4}n.$ (21)
In the following, we prove $T(n)=\mathcal{O}(n\log\log n)$ by mathematical induction.
First, for $n<\max(n_{2},n_{3},n_{4},\tau)\eqcolon n_{0}$, there exists a constant $C~{}(>0)$ such that $T(n)\leq C$. That is, for $n<n_{0}$, our Learned Sort terminates in a finite number of operations.
Next, assume that there exists a constant $C~{}(>0)$ and $l~{}(>0)$ such that $T(n)\leq C+ln\log\log n$ for all $n<k$, where $k$ is an integer such that $k\geq n_{0}$. Then, from $k\geq n_{2}$,
$\begin{split}T(k)&\leq R(k)+\mathbb{E}\left[\sum_{j=1}^{\gamma+1}\mathbbm{1}[|\bm{c}_{j}|\geq\lfloor k^{d}\rfloor]\cdot S(|\bm{c}_{j}|)+\mathbbm{1}[|\bm{c}_{j}|<\lfloor k^{d}\rfloor]\cdot T(|\bm{c}_{j}|)\right]\\ &\leq l_{2}k+\Pr\left[\exists j,|\bm{c}_{j}|\geq\lfloor k^{d}\rfloor\right]\cdot\mathbb{E}\left[\sum_{j=1}^{\gamma+1}S(|\bm{c}_{j}|)\right]+\mathbb{E}\left[\sum_{j=1}^{\gamma+1}\mathbbm{1}[|\bm{c}_{j}|<\lfloor k^{d}\rfloor]\cdot T(|\bm{c}_{j}|)\right]\\ &\leq l_{2}k+\Pr[\exists j,|\bm{c}_{j}|\geq\lfloor k^{d}\rfloor]\cdot\{C_{1}(\gamma+1)+l_{1}k\log k\}+\mathbb{E}\left[\sum_{i=1}^{\gamma+1}T(\min(\lfloor k^{d}\rfloor,|\bm{c}_{j}|))\right].\end{split}$ (22)
Here, from $k\geq n_{3}$ and $k\geq n_{4}$,
$\begin{split}\Pr[\exists j,|\bm{c}_{j}|\geq\lfloor k^{d}\rfloor]\cdot\{C_{1}(\gamma+1)+l_{1}k\log k\}&\leq\frac{l_{3}}{\log k}\cdot(C_{1}l_{4}k+l_{1}k\log k)\\ &\leq(C_{1}l_{3}l_{4}+l_{1}l_{3})k.\end{split}$ (23)
Also, from the assumption of induction and $k\geq n_{4}$,
$\begin{split}\mathbb{E}\left[\sum_{i=1}^{\gamma+1}T(\min(\lfloor k^{d}\rfloor,|\bm{c}_{j}|))\right]&\leq\mathbb{E}\left[\sum_{i=1}^{\gamma+1}\left\{C+l\cdot\min(\lfloor k^{d}\rfloor,|\bm{c}_{j}|)\log\log\min(\lfloor k^{d}\rfloor,|\bm{c}_{j}|)\right\}\right]\\ &\leq\mathbb{E}\left[C(\gamma+1)+\sum_{i=1}^{\gamma+1}l\cdot|\bm{c}_{j}|\log\log\lfloor k^{d}\rfloor\right]\\ &\leq Cl_{4}k+lk\log\log\lfloor k^{d}\rfloor\\ &\leq Cl_{4}k+lk\log d+lk\log\log k.\end{split}$ (24)
Therefore,
$\begin{split}T(k)&\leq l_{2}k+(C_{1}l_{3}l_{4}+l_{1}l_{3})k+Cl_{4}k+lk\log d+lk\log\log k\\ &\leq\left\{l_{2}+C_{1}l_{3}l_{4}+l_{1}l_{3}+Cl_{4}-l\log\frac{1}{d}\right\}k+(C+lk\log\log k).\end{split}$ (25)
Therefore, if we take $l$ such that
$\frac{l_{2}+C_{1}l_{3}l_{4}+l_{1}l_{3}+Cl_{4}}{\log\frac{1}{d}}\leq l,$ (26)
then $T(k)\leq C+lk\log\log k$ (note that the left side of Equation 26 is a constant independent of $k$).
Hence, by mathematical induction, it is proved that there exists a constant $C~{}(>0)$ and $l~{}(>0)$ such that $T(n)\leq C+ln\log\log n$ for all $n\in\mathbb{N}$. ∎
A.3 Proof of Lemma 3.4
To prove Lemma 3.4, we first present the following lemma.
Lemma A.1.
Let $\bm{e}~{}(\in\mathcal{I}^{n})$ be a sorted version of $\bm{x}_{\mathcal{I}}~{}(\in\mathcal{I}^{n})$ and $\Delta\coloneqq(x_{\mathrm{max}}-x_{\mathrm{min}})/\beta$ (where $x_{\mathrm{min}}$ and $x_{\mathrm{max}}$ are the minimum and maximum values of $\bm{x}_{\mathcal{I}}$, respectively).
Also, define the set $\mathcal{S}_{r},\mathcal{T}_{r}$ as follows ($r=1,\dots,n$):
$\mathcal{S}_{r}=\{k\mid e_{\max(1,r-\delta/2)}+\Delta<e_{k}\leq e_{r}-\Delta\},~{}~{}~{}~{}~{}~{}~{}~{}\mathcal{T}_{r}=\{k\mid e_{r}+\Delta\leq e_{k}<e_{\min(r+\delta/2,n)}-\Delta\}.$ (27)
Using this definition, define $Y_{jr},Z_{jr},Y_{r},Z_{r}$ as follows ($j=1,\dots,\alpha,~{}r=1,\dots,n$):
$Y_{jr}=\begin{cases}1&(j\in\mathcal{S}_{r})\\ 0&(\mathrm{else})\end{cases},~{}~{}~{}~{}~{}~{}~{}~{}Z_{jr}=\begin{cases}1&(j\in\mathcal{T}_{r})\\ 0&(\mathrm{else})\end{cases},$ (28)
$Y_{r}=\sum_{j=1}^{\alpha}Y_{jr},~{}~{}~{}~{}~{}~{}~{}~{}Z_{r}=\sum_{j=1}^{\alpha}Z_{jr}.$ (29)
If the size of the bucket to which $e_{r}$ is allocated is greater than or equal to $\delta$, then the following holds:
$\left(r\geq\frac{\delta}{2}+1~{}\land~{}Y_{r}\leq\left\lfloor\frac{\alpha}{\gamma}\right\rfloor\right)~{}\lor~{}\left(r\leq n-\frac{\delta}{2}~{}\land~{}Z_{r}\leq\left\lfloor\frac{\alpha}{\gamma}\right\rfloor\right).$ (30)
Proof.
We prove the contraposition of the lemma. That is, we prove that $e_{r}$ is allocated to a bucket smaller than $\delta$ under the assumption that $\left(r<\frac{\delta}{2}+1~{}\lor~{}Y_{r}>\left\lfloor\frac{\alpha}{\gamma}\right\rfloor\right)~{}\land~{}\left(r>n-\frac{\delta}{2}~{}\lor~{}Z_{r}>\left\lfloor\frac{\alpha}{\gamma}\right\rfloor\right)$.
For convenience, we hypothetically define $e_{0}=-\infty,e_{n+1}=\infty$, and assign $e_{0}$ to the 0th bucket and $e_{n+1}$ to the $(\gamma+2)$-th bucket. The size of the 0th bucket and the $(\gamma+2)$-th bucket are both 1.
First, we prove that $e_{r}$ and $e_{\min\left(r+\delta/2,n+1\right)}$ are assigned to different buckets. When $n-\delta/2<r\leq n$, $e_{r}$ and $e_{\min\left(r+\delta/2,n+1\right)}=e_{n+1}$ are obviously assigned to different buckets. When $r\leq n-\delta/2$, the ID of the bucket to which $e_{r}$ is assigned is
$\begin{split}\lfloor\tilde{F}(e_{r})\gamma\rfloor+1&=\left\lfloor\frac{\gamma}{\alpha}b_{i(e_{r})}\right\rfloor+1\\ &\leq\frac{\gamma}{\alpha}b_{i(e_{r})}+1\\ &=\frac{\gamma}{\alpha}\left|\{j\mid i(a_{j})\leq i(e_{r})\}\right|+1\\ &\leq\frac{\gamma}{\alpha}\left|\{j\mid a_{j}\leq e_{r}+\Delta\}\right|+1.\end{split}$ (31)
The ID of the bucket to which $e_{\min\left(r+\delta/2,n+1\right)}=e_{r+\delta/2}$ is assigned is
$\begin{split}\lfloor\tilde{F}(e_{r+\delta/2})\gamma\rfloor+1&=\left\lfloor\frac{\gamma}{\alpha}b_{i(e_{r+\delta/2})}\right\rfloor+1\\ &>\frac{\gamma}{\alpha}b_{i(e_{r+\delta/2})}\\ &=\frac{\gamma}{\alpha}\left|\{j\mid i(a_{j})\leq i(e_{r+\delta/2})\}\right|\\ &\geq\frac{\gamma}{\alpha}\left|\{j\mid a_{j}\leq e_{r+\delta/2}-\Delta\}\right|.\end{split}$ (32)
Thus, taking the difference between these two bucket IDs,
$\begin{split}\left(\lfloor\tilde{F}(e_{r+\delta/2})\gamma\rfloor+1\right)-\left(\lfloor\tilde{F}(e_{r})\gamma\rfloor+1\right)&>\frac{\gamma}{\alpha}\left|\{j\mid a_{j}\leq e_{r+\delta/2}-\Delta\}\right|-\left(\frac{\gamma}{\alpha}\left|\{j\mid a_{j}\leq e_{r}+\Delta\}\right|+1\right)\\ &=\frac{\gamma}{\alpha}\left|\{j\mid e_{r}+\Delta<a_{j}\leq e_{r+\delta/2}-\Delta\}\right|-1\\ &=\frac{\gamma}{\alpha}\left|\mathcal{T}_{r}\right|-1\\ &=\frac{\gamma}{\alpha}\sum_{j=1}^{\alpha}Z_{jr}-1\\ &=\frac{\gamma}{\alpha}Z_{r}-1\\ &\geq 0.\end{split}$ (33)
Therefore, $e_{r}$ and $e_{\min\left(r+\delta/2,n+1\right)}$ are assigned to different buckets.
In the same way, we can prove that $e_{\max(0,r-\delta/2)}$ and $e_{r}$ are also assigned to different buckets. Thus, the size of the bucket to which $e_{r}$ is assigned is at most $\delta-1$ (at most from $e_{\max(0,r-\delta/2)+1}$ to $e_{\min\left(r+\delta/2,n+1\right)-1}$), and the contraposition of the lemma is proved. ∎
Using Lemma A.1, we can prove Lemma 3.4.
Proof.
[…]
Thus, when $r\geq\frac{\delta}{2}+1$,
$\begin{split}\mathbb{E}\left[\frac{\delta}{2}-\left|\mathcal{S}_{r}\right|\right]&=\mathbb{E}\left[\frac{\delta}{2}-\left|\{k\mid e_{r-\delta/2}+\Delta<e_{k}\leq e_{r}-\Delta\}\right|\right]\\ &=\mathbb{E}\left[\left|\{k\mid e_{r-\delta/2}<e_{k}\leq e_{r-\delta/2}+\Delta\}\right|\right]+\mathbb{E}\left[\left|\{k\mid e_{r}-\Delta<e_{k}\leq e_{r}\}\right|\right]\\ &\leq nq+nq\\ &\leq\frac{2\sigma_{2}n}{\sigma_{1}\beta}.\end{split}$ (35)
Thus, when $r\geq\frac{\delta}{2}+1$,
$\begin{split}\mathbb{E}[Y_{r}]&=\frac{\alpha}{n}\mathbb{E}\left[\left|\mathcal{S}_{r}\right|\right]\\ &=\frac{\alpha}{n}\left(\frac{\delta}{2}-\mathbb{E}\left[\frac{\delta}{2}-\left|\mathcal{S}_{r}\right|\right]\right)\\ &\geq\frac{\alpha\delta}{2n}-\frac{2\sigma_{2}\alpha}{\sigma_{1}\beta}\\ &=\frac{\alpha K}{\gamma}.\end{split}$ (36)
Here, when $K\geq 1$, we have
$10\leq 1-\frac{\alpha}{\gamma\mathbb{E}[Y_{r}]}<$ 1 . (37)
Therefore, from the Chernoff bound,
$\begin{split}\Pr\left[Y_{r}\leq\frac{\alpha}{\gamma}\right]&=\Pr\left[Y_{r}\leq\left\{1-\left(1-\frac{\alpha}{\gamma\mathbb{E}[Y_{r}]}\right)\right\}\mathbb{E}\left[Y_{r}\right]\right]\\ &\leq\exp\left\{-\frac{1}{2}\left(1-\frac{\alpha}{\gamma\mathbb{E}[Y_{r}]}\right)^{2}\mathbb{E}\left[Y_{r}\right]\right\}\\ &\leq\exp\left\{-\frac{\alpha K}{2\gamma}\left(1-\frac{1}{K}\right)^{2}\right\}.\end{split}$ (38)
In the same way, we can prove that when $r\leq n-\frac{\delta}{2}$,
$\Pr\left[Z_{r}\leq\frac{\alpha}{\gamma}\right]\leq\exp\left\{-\frac{\alpha K}{2\gamma}\left(1-\frac{1}{K}\right)^{2}\right\}.$ (39)
Thus, by defining $E_{r}$ to be the event “$e_{r}$ is allocated to a bucket with size greater than or equal to $\delta$,” from Lemma A.1,
$\begin{split}\Pr[E_{r}]&\leq\Pr\left[\left(r\geq\frac{\delta}{2}+1~{}\land~{}Y_{r}\leq\left\lfloor\frac{\alpha}{\gamma}\right\rfloor\right)~{}\lor~{}\left(r\leq n-\frac{\delta}{2}~{}\land~{}Z_{r}\leq\left\lfloor\frac{\alpha}{\gamma}\right\rfloor\right)\right]\\ &\leq\Pr\left[\left(r\geq\frac{\delta}{2}+1~{}\land~{}Y_{r}\leq\left\lfloor\frac{\alpha}{\gamma}\right\rfloor\right)\right]+\Pr\left[\left(r\leq n-\frac{\delta}{2}~{}\land~{}Z_{r}\leq\left\lfloor\frac{\alpha}{\gamma}\right\rfloor\right)\right]\\ &\leq\begin{cases}\exp\left\{-\frac{\alpha K}{2\gamma}\left(1-\frac{1}{K}\right)^{2}\right\},&\left(r<\frac{\delta}{2}+1~{}\lor~{}r>n-\frac{\delta}{2}\right)\\ 2\exp\left\{-\frac{\alpha K}{2\gamma}\left(1-\frac{1}{K}\right)^{2}\right\},&(\mathrm{else})\end{cases}\\ &\leq 2\exp\left\{-\frac{\alpha K}{2\gamma}\left(1-\frac{1}{K}\right)^{2}\right\}.\end{split}$ (40)
Therefore,
$\mathbb{E}\left[\sum_{r=1}^{n}\mathbbm{1}[E_{r}]\right]\leq 2n\exp\left\{-\frac{\alpha K}{2\gamma}\left(1-\frac{1}{K}\right)^{2}\right\}$ (41)
Then, noting that the number of buckets with size greater than or equal to $\delta$ is less than or equal to $\sum_{r=1}^{n}\mathbbm{1}[E_{r}]/\delta$,
$\mathbb{E}\left[|\{j\mid|\bm{c}_{j}|>\delta\}|\right]\leq\frac{2n}{\delta}\exp\left\{-\frac{\alpha K}{2\gamma}\left(1-\frac{1}{K}\right)^{2}\right\}.$ (42)
Then, from Markov’s inequality, we have
$\begin{split}\Pr[\exists j,|\bm{c}_{j}|>\delta]&=\Pr[|\{j\mid|\bm{c}_{j}|>\delta\}|\geq 1]\\ &\leq\frac{2n}{\delta}\exp\left\{-\frac{\alpha K}{2\gamma}\left(1-\frac{1}{K}\right)^{2}\right\}.\end{split}$ (43)